# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ C.2 Matching finite-sample error bound
$\leq\|\psi_{\pi}\|_{\Lambda^{-1}}\sqrt{\frac{2\log(8d/\delta)}{T}}\left(1+\frac{K_{\delta/4}}{2}\right)+\frac{\|\psi_{\pi}\|_{\Lambda^{-1}}}{T}4\sqrt{2dL}\log\left(16(d+1)/\delta\right)\left(1+\frac{K_{\delta/8}}{2}\right)$
$+\frac{\|\psi_{\pi}\|_{\Lambda^{-1}}LL_{\theta}\log(8d/\delta)}{T}\left(1+4\sqrt{\frac{\log(8d/\delta)L}{T}}\right)$
$\leq\|\psi_{\pi}\|_{\Lambda^{-1}}\left(\sqrt{\frac{2\log(8d/\delta)}{T}}+\frac{4\sqrt{2dL}}{T}\log\left(16(d+1)/\delta\right)\right)\left(1+\sqrt{\frac{\log\left(16(d+1)/\delta\right)}{2T}}+\frac{\log(16(d+1)/\delta)L}{3T}\right)$
$+\frac{\|\psi_{\pi}\|_{\Lambda^{-1}}LL_{\theta}\log(8d/\delta)}{T}\left(1+4\sqrt{\frac{\log(8d/\delta)L}{T}}\right)$
$\leq\frac{1}{\sqrt{T}}E_{1}+\|\psi_{\pi}\|_{\Lambda^{-1}}\frac{4\sqrt{2dL}}{T}\log\left(16(d+1)/\delta\right)+\frac{LL_{\theta}\log(8d/\delta)}{T}\|\psi_{\pi}\|_{\Lambda^{-1}}\left(1+4\sqrt{\frac{\log(8d/\delta)L}{T}}\right)$
$+\|\psi_{\pi}\|_{\Lambda^{-1}}\left(\sqrt{\frac{2\log\left(16(d+1)/\delta\right)}{T}}+\frac{4\sqrt{2dL}}{T}\log\left(16(d+1)/\delta\right)\right)\left(\sqrt{\frac{\log\left(16(d+1)/\delta\right)}{2T}}+\frac{\log(16(d+1)/\delta)L}{3T}\right)$
@@ C.3 Additional Lemmas
We first consider the case $\nu_{i}\in\operatorname{Im}\left(\Lambda_{i}\right)$ for every $i\in[n]$. In this case, $\lambda\Lambda_{i}^{\dagger}\nu_{i}$ is a solution of $\Lambda_{i}\mu_{i}=\lambda\nu_{i}$ for any $\lambda.$ Thus, for $\lambda^{*}=\frac{2\varepsilon}{\sum_{i=1}^{n}\nu_{i}^{\top}\Lambda_{i}^{\dagger}\nu_{i}}$ and $\mu_{i}^{*}=\lambda_{*}\Lambda_{i}^{\dagger}\nu_{i}$, we have $\Lambda_{i}\mu_{i}^{*}=\lambda^{*}\nu_{i}$ and
@@
When $\nu_{j}\notin\operatorname{Im}\left(\Lambda_{j}\right)$ for some $j\in[n]$, $\Lambda_{j}\mu_{j}=\lambda\nu_{j}$ implies $\lambda=0$. Thus, any optimal point $\mu^{*}$ must verify $\Lambda_{i}\mu_{i}^{*}=0$ for every $i\in[n]$ which implies that $\frac{1}{2}\sum_{i=1}^{n}{\mu_{i}^{*}}^{\top}\Lambda_{i}\mu_{i}^{*}=0$. ∎
@@
Let $\left(e_{t}\right)_{t=1,\dots,T}$ be a martingale difference sequence with respect to a filtration $(\mathcal{F}_{t})_{t=0,\dots,T-1}$ (i.e $\mathbb{E}\left[e_{t}|\mathcal{F}_{t-1}\right]=0$ for all $t$), and let $V_{t}=\operatorname{Var}\left(e_{t}|\mathcal{F}_{t-1}\right)$. If for all $c\geq 0$ and all $t=1,\dots,T$, we have $\mathbb{E}\left[e^{ce_{t}}|\mathcal{F}_{t-1}\right]\leq e^{c^{2}V_{t}/2}$, then for all $\varepsilon,\sigma>0$,
$\mathbb{P}\left(\left|\sum_{t=1}^{T}e_{t}\right|\geq\varepsilon\text{ and }\sum_{t=1}^{T}V_{t}\leq\sigma^{2}\right)\leq 2\exp\left(-\frac{\varepsilon^{2}}{2\sigma^{2}}\right).$
@@
$\mathbb{P}\left(\sum_{t=1}^{T}e_{t}\geq\varepsilon\text{ and }\sum_{t=1}^{T}V_{t}\leq\sigma^{2}\right)\leq\exp\left(-\frac{\varepsilon^{2}}{2\sigma^{2}}\right).$
@@
Let $\left(W_{t}\right)_{t=1,\dots,T}$ be a conditionally symmetric martingale difference sequence of $\mathbb{R}^{d\times d^{\prime}}$ with respect to a filtration $(\mathcal{F}_{t})_{t=0,\dots,T-1}$, and let $V_{1,t}\in\mathbb{R}^{d\times d}$, $V_{2,t}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ be $\mathcal{F}_{t-1}$-measurable symmetric matrices. Denote by $Z_{1,t}\in\mathbb{R}^{d\times d}$ and $Z_{2,t}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ any symmetric matrices such that $Z_{1,t}^{2}=W_{t}W_{t}^{\top}$ and $Z_{2,t}^{2}=W_{t}^{\top}W_{t}$. If for some nonnegative function $g$, for all $c\in\mathbb{R}$ and all $t=1,\dots,T$, $\mathbb{E}\left[e^{cZ_{1,t}}|\mathcal{F}_{t-1}\right]\preccurlyeq e^{g(c)V_{1,t}}$ and $\mathbb{E}\left[e^{cZ_{2,t}}|\mathcal{F}_{t-1}\right]\preccurlyeq e^{g(c)V_{2,t}}$, then for all $\varepsilon,\sigma>0$,
$\mathbb{P}\left(\lambda_{\max}\left(\sum_{t=1}^{T}W_{t}\right)\geq\varepsilon\text{ and }\max\left(\lambda_{\max}\left(\sum_{t=1}^{T}V_{1,t}\right),\lambda_{\max}\left(\sum_{t=1}^{T}V_{2,t}\right)\right)\leq\sigma^{2}\right)\leq(d+d^{\prime})\inf_{c\in\mathbb{R}}\exp\left(-c\varepsilon+g(c)\sigma^{2}\right).$
@@
Let $Y_{t}:=\begin{pmatrix}0&W_{t}\\ W_{t}^{\top}&0\end{pmatrix}$ and $V_{t}:=\begin{pmatrix}V_{1,t}&0\\ 0&V_{2,t}\end{pmatrix}$. We will use Theorem 2.3 of (Tropp, 2011a) on the martingale difference sequence $(Y_{t})_{t\geq 0}$ of $\mathbb{R}^{(d+d^{\prime})^{2}}$. To do so, we must prove that for all $c\in\mathbb{R}$ , $\mathbb{E}\left[e^{cY_{t}}|\mathbb{F}_{t-1}\right]\preccurlyeq e^{g(c)V_{t}}.$ It is straightforward to check that for all $k\geq 1$,
@@
$\mathbb{E}\left[\operatorname{cosh}\left(cZ_{1,t}\right)|\mathcal{F}_{t-1}\right]\preccurlyeq e^{g(c)V_{2,t}}.$
@@
$\mathbb{E}\left[e^{cY_{t}}|\mathcal{F}_{t-1}\right]$ $\preccurlyeq\begin{pmatrix}e^{g(c)V_{1,t}}&0\\ 0&e^{g(c)V_{2,t}}\end{pmatrix}$
@@
$\mathbb{P}\left(\lambda_{\max}\left(\sum_{t=1}^{T}Y_{t}\right)\geq\varepsilon\text{ and }\lambda_{\max}\left(\sum_{t=1}^{T}V_{t}\right)\leq\sigma^{2}\right)\leq(d+d^{\prime})\inf_{c\in\mathbb{R}}\exp\left(-c\varepsilon+g(c)\sigma^{2}\right).$
Notice that
$\lambda_{\max}\left(\sum_{t=1}^{T}Y_{t}\right)=\lambda_{\max}\left(\sum_{t=1}^{T}W_{t}\right)\text{ and }\lambda_{\max}\left(\sum_{t=1}^{T}V_{t}\right)=\max\left(\lambda_{\max}\left(\sum_{t=1}^{T}V_{1,t}\right),\lambda_{\max}\left(\sum_{t=1}^{T}V_{2,t}\right)\right),$
@@
Let $\left(W_{t}\right)_{t=1,\dots,T}$ be a conditionally symmetric martingale difference sequence of $\mathbb{R}^{d\times d^{\prime}}$ with respect to a filtration $(\mathcal{F}_{t})_{t=0,\dots,T-1}$, and let $V_{1,t}:=\mathbb{E}\left[W_{t}W_{t}^{\top}|\mathcal{F}_{t-1}\right]$, $V_{2,t}:=\mathbb{E}\left[W_{t}^{\top}W_{t}|\mathcal{F}_{t-1}\right]$. Denote by $Z_{1,t}\in\mathbb{R}^{d\times d}$ and $Z_{2,t}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ any symmetric matrices such that $Z_{1,t}^{2}=W_{t}W_{t}^{\top}$ and $Z_{2,t}^{2}=W_{t}^{\top}W_{t}$. If for all $c\in\mathbb{R}$ and all $t=1,\dots,T$, $\mathbb{E}\left[e^{cZ_{1,t}}|\mathcal{F}_{t-1}\right]\preccurlyeq e^{c^{2}V_{1,t}/2}$ and $\mathbb{E}\left[e^{cZ_{2,t}}|\mathcal{F}_{t-1}\right]\preccurlyeq e^{c^{2}V_{2,t}/2}$, then for all $\varepsilon,\sigma>0$,
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}W_{t}\right\|_{{\textup{op}}}\geq\varepsilon\ \mathrm{and}\ \max\left(\left\|\sum_{t=1}^{T}V_{1,t}\right\|_{{\textup{op}}},\left\|\sum_{t=1}^{T}V_{2,t}\right\|_{{\textup{op}}}\right)\leq\sigma^{2}\right)\leq 2(d+d^{\prime})\exp\left(-\frac{\varepsilon^{2}}{2\sigma^{2}}\right).$
@@
We choose $g(c)=c^{2}/2$ in the statement of Proposition C.10. Note that $\inf_{c\in\mathbb{R}}\exp\left(-c\varepsilon+g(c)\sigma^{2}\right)=\exp\left(-\frac{\varepsilon^{2}}{2\sigma^{2}}\right)$, and that $V_{1,t}$ and $V_{2,t}$ are positive semi-definite matrices, which implies that $\lambda_{\max}\left(\sum_{t=1}^{T}V_{1,t}\right)=\|\sum_{t=1}^{T}V_{1,t}\|_{{\textup{op}}}$ and $\lambda_{\max}\left(\sum_{t=1}^{T}V_{2,t}\right)=\|\sum_{t=1}^{T}V_{2,t}\|_{{\textup{op}}}$. The same bound applies to the conditionally symmetric martingale difference sequence $\left(-W_{t}\right)$, and a union bound concludes the proof. ∎
@@
We first note that the max-norm guarantee [^11] The max-norm is denoted by $\|.\|_{\infty}$ in (Stojanovic et al., 2023). of Theorem 1 of (Stojanovic et al., 2023) can be straightforwardly generalized to the case where the samples are not necessarily selected uniformly at random by leveraging the two-to-infinity norm guarantees of Proposition 4.1. We state the general result below.
@@
Let $\delta\in\left(0,1\right)$. If $T\geq\frac{c\mu^{4}\kappa^{2}r^{2}\left(m+n\right)}{\omega_{\min}mn}\log^{3}\left(e(m+n)T/\delta\right)$ for some universal constant $c>0$, then for some universal constant $C>0$, with probability larger than $1-\delta$,
$\|M-\widehat{M}\|_{\max}\leq C\mu^{11/2}\kappa^{2}r^{3/2}\|M\|_{\max}\frac{(m+n)^{3/2}}{\sqrt{T\omega_{\min}mn}\left(m\wedge n\right)}\log^{3/2}\left(\frac{(m+n)T}{\delta}\right).$
@@
When the context distribution and behavior policy are not uniform, the key term $\mathcal{B}$ in the statement of Theorem 1 of (Stojanovic et al., 2023) becomes
$\mathcal{B}=\frac{1}{\sqrt{T\omega_{\min}}}\left(\sqrt{(m+n)\log\left(\frac{e(m+n)T}{\delta}\right)}+\log^{3/2}\left(\frac{e(m+n)T}{\delta}\right)\right).$
By following the arguments in the Proof of Theorem 1 of (Stojanovic et al., 2023), we obtain that there is a universal constant $c>0$ such that as soon as $T\geq\frac{c\mu^{4}\kappa^{2}r^{2}\left(m+n\right)}{\omega_{\min}mn}\log^{3}\left(e(m+n)T/\delta\right)$, then for some universal constant $C>0$, with probability larger than $1-\delta$,
$\|M-\widehat{M}\|_{\max}\leq C\left(\mu^{11/2}\kappa^{2}r^{1/2}+\mu^{3}\kappa r^{3/2}\frac{m+n}{\sqrt{mn}}\right)\frac{\|M\|_{\max}\mathcal{B}}{m\wedge n}.$
The final result follows from $\mu,\kappa,r\geq 1$. ∎
@@
$T\gtrsim\frac{\mu^{11}\kappa^{4}r^{3}\|M\|_{\max}^{2}(m+n)^{3}}{\omega_{\min}mn(m\wedge n)^{2}\varepsilon^{2}}\log^{3}\!\left(\frac{(m+n)T}{\delta}\right)\!$
@@
$T\geq\frac{c\mu^{4}\kappa^{2}r^{2}\left(m+n\right)}{\omega_{\min}mn}\log^{3}\left(\frac{(m+n)T}{\delta}\right).$
Finally, note that when $\|M\|_{\max}\gtrsim 1,$ the second condition is weaker and can consequently be dropped.
@@
Recall that $M_{i,j}=\phi_{i,j}^{\top}\theta+\epsilon_{i,j}.$ In this section, the following additional notations will be used:
$\Lambda_{\phi}=\sum_{i,j}\omega_{i,j}\phi_{i,j}\phi_{i,j}^{\top},\widehat{\Lambda}_{\tau}=\frac{1}{T-T_{1}}\left(\sum_{t=T_{1}+1}^{T}\phi_{i_{t},j_{t}}\phi_{i_{t},j_{t}}^{\top}+\tau I_{d}\right)\text{ and }\phi_{\pi}=\sum_{i,j}\omega^{\pi}_{i,j}\phi_{i,j}.$
We also define
$\epsilon:=\max_{i\in[m],j\in[n]}\epsilon_{i,j}$
as the misspecification error of the misspecified linear bandit problem. Finally, to ease the notation, we do not write the dependency in $\pi$ of the estimator $\hat{v}^{\pi}_{\text{\tt RS-PE}}$. We summarize RS-PE in Algorithm 3.

### v2

@@ C.2 Matching finite-sample error bound
$\leq\|\psi_{\pi}\|_{\Lambda^{-1}}\sqrt{\frac{2\sigma^{2}C_{1}}{T}}\left(1+\frac{K_{\delta/4}}{2}\right)+\frac{\|\psi_{\pi}\|_{\Lambda^{-1}}}{T}4\sqrt{2d\sigma^{2}L_{\psi}}C_{2}\left(1+\frac{K_{\delta/8}}{2}\right)$
$+\frac{\|\psi_{\pi}\|_{\Lambda^{-1}}L_{\psi}L_{\theta}C_{1}}{T}\left(1+4\sqrt{\frac{C_{1}L_{\psi}}{T}}\right)$
$\leq\|\psi_{\pi}\|_{\Lambda^{-1}}\left(\sqrt{\frac{2\sigma^{2}C_{1}}{T}}+\frac{4\sqrt{2d\sigma^{2}L_{\psi}}}{T}C_{2}\right)\left(1+\sqrt{\frac{C_{2}L_{\psi}}{2T}}+\frac{C_{2}L_{\psi}}{3T}\right)$
$+\frac{\|\psi_{\pi}\|_{\Lambda^{-1}}L_{\psi}L_{\theta}C_{1}}{T}\left(1+4\sqrt{\frac{C_{1}L_{\psi}}{T}}\right)$
$\leq\frac{1}{\sqrt{T}}E_{1}+\|\psi_{\pi}\|_{\Lambda^{-1}}\frac{4\sqrt{2d\sigma^{2}L_{\psi}}}{T}C_{2}+\frac{L_{\psi}L_{\theta}C_{1}}{T}\|\psi_{\pi}\|_{\Lambda^{-1}}\left(1+4\sqrt{\frac{C_{1}L_{\psi}}{T}}\right)$
$+\|\psi_{\pi}\|_{\Lambda^{-1}}\left(\sqrt{\frac{2\sigma^{2}C_{2}}{T}}+\frac{4\sqrt{2d\sigma^{2}L_{\psi}}}{T}C_{2}\right)\left(\sqrt{\frac{C_{2}L_{\psi}}{2T}}+\frac{C_{2}L_{\psi}}{3T}\right)$
@@ C.3 Additional Lemmas
  • If $\nu_{i}\in\operatorname{Im}\left(\Lambda_{i}\right)$ for every $i\in[n]$, $\lambda\Lambda_{i}^{\dagger}\nu_{i}$ is a solution of $\Lambda_{i}\mu_{i}=\lambda\nu_{i}$ for any $\lambda.$ Thus, for $\lambda^{*}=\frac{2\varepsilon}{\sum_{i=1}^{n}\nu_{i}^{\top}\Lambda_{i}^{\dagger}\nu_{i}}$ and $\mu_{i}^{*}=\lambda_{*}\Lambda_{i}^{\dagger}\nu_{i}$, we have $\Lambda_{i}\mu_{i}^{*}=\lambda^{*}\nu_{i}$ and
@@
  • If $\nu_{j}\notin\operatorname{Im}\left(\Lambda_{j}\right)$ for some $j\in[n]$, $\Lambda_{j}\mu_{j}=\lambda\nu_{j}$ implies $\lambda=0$. Thus, any optimal point $\mu^{*}$ must verify $\Lambda_{i}\mu_{i}^{*}=0$ for every $i\in[n]$ which implies that $\frac{1}{2}\sum_{i=1}^{n}{\mu_{i}^{*}}^{\top}\Lambda_{i}\mu_{i}^{*}=0$.
@@
Let $\left(e_{t}\right)_{t=1,\dots,T}$ be a martingale difference sequence with respect to a filtration $(\mathcal{F}_{t})_{t=0,\dots,T-1}$ (i.e., $\mathbb{E}\left[e_{t}|\mathcal{F}_{t-1}\right]=0$ for every $t$), and let $V_{t}=\operatorname{Var}\left(e_{t}|\mathcal{F}_{t-1}\right)$. If for all $c\geq 0$ and all $t=1,\dots,T$, we have $\mathbb{E}\left[e^{ce_{t}}|\mathcal{F}_{t-1}\right]\leq e^{c^{2}V_{t}/2}$, then for all $\varepsilon,\gamma>0$,
$\mathbb{P}\left(\left|\sum_{t=1}^{T}e_{t}\right|\geq\varepsilon\text{ and }\sum_{t=1}^{T}V_{t}\leq\gamma^{2}\right)\leq 2\exp\left(-\frac{\varepsilon^{2}}{2\gamma^{2}}\right).$
@@
$\mathbb{P}\left(\sum_{t=1}^{T}e_{t}\geq\varepsilon\text{ and }\sum_{t=1}^{T}V_{t}\leq\gamma^{2}\right)\leq\exp\left(-\frac{\varepsilon^{2}}{2\gamma^{2}}\right).$
@@
Let $\left(W_{t}\right)_{t=1,\dots,T}$ be a conditionally symmetric [^11] This means that $W_{t}$ and $-W_{t}$ have the same conditional distribution with respect to $\mathcal{F}_{t-1}$ for every $t$. martingale difference sequence of $\mathbb{R}^{d\times d^{\prime}}$ with respect to a filtration $(\mathcal{F}_{t})_{t=0,\dots,T-1}$, and let $V_{1,t}\in\mathbb{R}^{d\times d}$, $V_{2,t}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ be $\mathcal{F}_{t-1}$-measurable symmetric matrices. Denote by $Z_{1,t}\in\mathbb{R}^{d\times d}$ and $Z_{2,t}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ any symmetric matrices such that $Z_{1,t}^{2}=W_{t}W_{t}^{\top}$ and $Z_{2,t}^{2}=W_{t}^{\top}W_{t}$. If for some nonnegative function $g$, for all $c\in\mathbb{R}$ and all $t=1,\dots,T$, $\mathbb{E}\left[e^{cZ_{1,t}}|\mathcal{F}_{t-1}\right]\preceq e^{g(c)V_{1,t}}$ [^12] Recall that $N\preceq M$ means that $M-N$ is positive semi-definite. and $\mathbb{E}\left[e^{cZ_{2,t}}|\mathcal{F}_{t-1}\right]\preceq e^{g(c)V_{2,t}}$, then for all $\varepsilon,\gamma>0$,
$\mathbb{P}\left(\lambda_{\max}\left(\sum_{t=1}^{T}W_{t}\right)\geq\varepsilon\text{ and }\max\left(\lambda_{\max}\left(\sum_{t=1}^{T}V_{1,t}\right),\lambda_{\max}\left(\sum_{t=1}^{T}V_{2,t}\right)\right)\leq\gamma^{2}\right)\leq(d+d^{\prime})\inf_{c\in\mathbb{R}}\exp\left(-c\varepsilon+g(c)\gamma^{2}\right).$
@@
Let $Y_{t}:=\begin{pmatrix}0&W_{t}\\ W_{t}^{\top}&0\end{pmatrix}$ and $V_{t}:=\begin{pmatrix}V_{1,t}&0\\ 0&V_{2,t}\end{pmatrix}$. We will use Theorem 2.3 of (Tropp, 2011a) on the martingale difference sequence $(Y_{t})_{t\geq 0}$ of $\mathbb{R}^{(d+d^{\prime})^{2}}$. To do so, we must prove that for all $c\in\mathbb{R}$ , $\mathbb{E}\left[e^{cY_{t}}|\mathbb{F}_{t-1}\right]\preceq e^{g(c)V_{t}}.$ It is straightforward to check that for all $k\geq 1$,
@@
$\mathbb{E}\left[\operatorname{cosh}\left(cZ_{1,t}\right)|\mathcal{F}_{t-1}\right]\preceq e^{g(c)V_{2,t}}.$
@@
$\mathbb{E}\left[e^{cY_{t}}|\mathcal{F}_{t-1}\right]$ $\preceq\begin{pmatrix}e^{g(c)V_{1,t}}&0\\ 0&e^{g(c)V_{2,t}}\end{pmatrix}$
@@
$\mathbb{P}\left(\lambda_{\max}\left(\sum_{t=1}^{T}Y_{t}\right)\geq\varepsilon\text{ and }\lambda_{\max}\left(\sum_{t=1}^{T}V_{t}\right)\leq\gamma^{2}\right)\leq(d+d^{\prime})\inf_{c\in\mathbb{R}}\exp\left(-c\varepsilon+g(c)\gamma^{2}\right).$
Note that
$\lambda_{\max}\left(\sum_{t=1}^{T}Y_{t}\right)=\lambda_{\max}\left(\sum_{t=1}^{T}W_{t}\right)\quad\text{and}\quad\lambda_{\max}\left(\sum_{t=1}^{T}V_{t}\right)=\max\left(\lambda_{\max}\left(\sum_{t=1}^{T}V_{1,t}\right),\lambda_{\max}\left(\sum_{t=1}^{T}V_{2,t}\right)\right),$
@@
Let $\left(W_{t}\right)_{t=1,\dots,T}$ be a conditionally symmetric martingale difference sequence of $\mathbb{R}^{d\times d^{\prime}}$ with respect to a filtration $(\mathcal{F}_{t})_{t=0,\dots,T-1}$, and let $V_{1,t}:=\mathbb{E}\left[W_{t}W_{t}^{\top}|\mathcal{F}_{t-1}\right]$, $V_{2,t}:=\mathbb{E}\left[W_{t}^{\top}W_{t}|\mathcal{F}_{t-1}\right]$. Denote by $Z_{1,t}\in\mathbb{R}^{d\times d}$ and $Z_{2,t}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ any symmetric matrices such that $Z_{1,t}^{2}=W_{t}W_{t}^{\top}$ and $Z_{2,t}^{2}=W_{t}^{\top}W_{t}$. If for all $c\in\mathbb{R}$ and all $t=1,\dots,T$, $\mathbb{E}\left[e^{cZ_{1,t}}|\mathcal{F}_{t-1}\right]\preceq e^{c^{2}V_{1,t}/2}$ and $\mathbb{E}\left[e^{cZ_{2,t}}|\mathcal{F}_{t-1}\right]\preceq e^{c^{2}V_{2,t}/2}$, then for all $\varepsilon,\gamma>0$,
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}W_{t}\right\|_{{\textup{op}}}\geq\varepsilon\ \mathrm{and}\ \max\left(\left\|\sum_{t=1}^{T}V_{1,t}\right\|_{{\textup{op}}},\left\|\sum_{t=1}^{T}V_{2,t}\right\|_{{\textup{op}}}\right)\leq\gamma^{2}\right)\leq 2(d+d^{\prime})\exp\left(-\frac{\varepsilon^{2}}{2\gamma^{2}}\right).$
@@
We choose $g(c)=c^{2}/2$ in the statement of Proposition C.10. Note that $\inf_{c\in\mathbb{R}}\exp\left(-c\varepsilon+g(c)\gamma^{2}\right)=\exp\left(-\frac{\varepsilon^{2}}{2\gamma^{2}}\right)$, and that $V_{1,t}$ and $V_{2,t}$ are positive semi-definite matrices, which implies that $\lambda_{\max}\left(\sum_{t=1}^{T}V_{1,t}\right)=\|\sum_{t=1}^{T}V_{1,t}\|_{{\textup{op}}}$ and $\lambda_{\max}\left(\sum_{t=1}^{T}V_{2,t}\right)=\|\sum_{t=1}^{T}V_{2,t}\|_{{\textup{op}}}$. The same bound applies to the conditionally symmetric martingale difference sequence $\left(-W_{t}\right)$, and a union bound concludes the proof. ∎
@@
We first note that the max-norm guarantee [^13] The max-norm is denoted by $\|.\|_{\infty}$ in (Stojanovic et al., 2024). of Theorem 1 of (Stojanovic et al., 2024) can be generalized to the case where the samples are not necessarily selected uniformly at random by leveraging the two-to-infinity norm guarantees of Theorem 4.1. We state the general result below, where we recall that $L=\|M\|_{\max}\vee\sigma$ (note that the dependency in $\mu$ is also improved compared to Theorem 1 of (Stojanovic et al., 2024)).
@@
Let $\delta\in\left(0,1\right)$. If $T\geq\frac{cL^{2}\kappa^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$ for some universal constant $c>0$, then with probability larger than $1-\delta$,
$\|M-\widehat{M}\|_{\max}\lesssim\sqrt{\frac{L^{2}\mu^{6}\kappa^{4}r^{3}(m+n)}{T\omega_{\min}(m\wedge n)^{2}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}.$
@@
The proof is divided in two steps: first, we leverage the bound on $\epsilon_{\textup{Sub-Rec}}$ of Theorem 4.1 to bound $\|M-\widehat{M}\|_{2\to\infty},$ and then we bound $\|M-\widehat{M}\|_{\max}$ by relating it to $\epsilon_{\textup{Sub-Rec}}$ and $\|M-\widehat{M}\|_{2\to\infty}.$
Bounding $\|M-\widehat{M}\|_{2\to\infty}.$
Let $E:=M-\widetilde{M}$. From (45) in (Stojanovic et al., 2024),
$\|M-\widehat{M}\|_{2\to\infty}\leq\epsilon_{\textup{Sub-Rec}}\left(\|E\|_{{\textup{op}}}+\|M\|_{{\textup{op}}}\right)+\|U\|_{2\to\infty}\|E\|_{{\textup{op}}}.$
Hence [^14] Union bounds are performed throughout the proof. Since the relevant inequalities are stated up to a constant, they are not impacted., with probability larger than $1-\delta$, from Proposition B.5, and Theorem 4.1,
$\|M-\widehat{M}\|_{2\to\infty}$ $\lesssim\sigma_{1}\epsilon_{\textup{Sub-Rec}}+\mu\sqrt{\frac{r}{m}}\|E\|_{{\textup{op}}}$
$\lesssim\kappa\sqrt{\frac{L^{2}\mu^{2}\kappa^{2}r(m+n)}{T\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}$
$+\mu\sqrt{\frac{r}{m}}L\sqrt{\frac{m+n}{T\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}$
$\lesssim\sqrt{\frac{L^{2}\mu^{2}\kappa^{4}r(m+n)}{T\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}$
under $T\gtrsim\frac{L^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$ (note that by Lemma B.8, this condition is stronger than the condition of Proposition B.5).
Bounding $\|M-\widehat{M}\|_{\max}.$
We similarly write (see (Stojanovic et al., 2024)),
$\|M-\widehat{M}\|_{\max}\leq\epsilon_{\textup{Sub-Rec}}(\|M-\widehat{M}\|_{2\to\infty}+\|M\|_{2\to\infty})+\|V\|_{2\to\infty}\|M-\widehat{M}\|_{2\to\infty}.$
Note that by Lemma B.8, $\|M\|_{2\to\infty}\leq\sqrt{m}\|M\|_{\max}\leq\sigma_{1}\frac{\mu^{2}r}{\sqrt{n}}$. Furthermore, under
$T\gtrsim\frac{L^{2}\kappa^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right),$ (27)
we know that by Theorem 4.1, $\epsilon_{\textup{Sub-Rec}}\lesssim\frac{\mu\sqrt{r}}{\sqrt{m\wedge n}}$ with probability larger than $1-\delta$. Thus, with probability larger than $1-\delta$, under (27),
$\|\widehat{M}-M\|_{\max}$ $\lesssim\left(\epsilon_{\textup{Sub-Rec}}+\mu\sqrt{\frac{r}{n}}\right)\|M-\widehat{M}\|_{2\to\infty}+\sigma_{1}\epsilon_{\textup{Sub-Rec}}\frac{\mu^{2}r}{\sqrt{n}}$
$\lesssim\mu\sqrt{\frac{r}{m\wedge n}}\sqrt{\frac{L^{2}\mu^{2}\kappa^{4}r(m+n)}{T\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}$
$+\frac{\sigma_{1}}{\sigma_{r}}\frac{\mu^{2}r}{\sqrt{n}}\sqrt{\frac{L^{2}\mu^{2}\kappa^{2}r(m+n)}{T\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}$
$\lesssim\sqrt{\frac{L^{2}\mu^{6}\kappa^{4}r^{3}(m+n)}{T\omega_{\min}(m\wedge n)^{2}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}.$
@@
$T\gtrsim\frac{L^{2}\mu^{6}\kappa^{4}r^{3}(m+n)}{\omega_{\min}(m\wedge n)^{2}\varepsilon^{2}}\log^{3}\!\left(\frac{(m+n)T}{\delta}\right)\!$
@@
$T\gtrsim\frac{L^{2}\kappa^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right).$
By Lemma B.8, $\frac{\kappa^{2}}{\sigma_{r}^{2}}\leq\frac{\mu^{4}\kappa^{2}r^{2}}{mn\|M\|_{\max}^{2}}\leq\frac{\mu^{6}\kappa^{4}r^{3}}{(m\wedge n)^{2}\|M\|_{\max}^{2}}$. Since $\varepsilon\leq\|M\|_{\max}$, the second condition is weaker and can be dropped. ∎
@@
Recall that $M_{i,j}=\phi_{i,j}^{\top}\theta+\epsilon_{i,j}.$ In this appendix, the following additional notations will be used:
$\Lambda=\sum_{i,j}\omega_{i,j}\psi_{i,j}\psi_{i,j}^{\top},\Lambda_{\phi}=\sum_{i,j}\omega_{i,j}\phi_{i,j}\phi_{i,j}^{\top},\widehat{\Lambda}_{\tau}=\frac{1}{T-T_{1}}\left(\sum_{t=T_{1}+1}^{T}\phi_{i_{t},j_{t}}\phi_{i_{t},j_{t}}^{\top}+\tau I_{d}\right)\text{ and }\phi_{\pi}=\sum_{i,j}\omega^{\pi}_{i,j}\phi_{i,j}.$
We also recall that
$\epsilon_{\max}=\max_{i\in[m],j\in[n]}|\epsilon_{i,j}|$
is the misspecification of the misspecified linear bandit, and that the latter is of dimension $d=r(m+n)-r^{2}$. Finally, to ease the notation, we do not write the dependency in $\pi$ of the estimator $\hat{v}^{\pi}_{\text{\tt RS-PE}}$. We summarize RS-PE in Algorithm 3.